OpenStudy (cmtboy2):

Evaluate the expression, using the given value of the variable. 2 – 8x + 3x when x = 5 A. –53 B. –23 C. –15 D.57

Nnesha (nnesha):

substitute x for its value

OpenStudy (cmtboy2):

the value for x is 5

Nnesha (nnesha):

right so replace x with 5

OpenStudy (cmtboy2):

i did

Nnesha (nnesha):

show me what did yo uget after replacing x with 5

OpenStudy (dakid88):

2 - 8x + 3x x = 5 2 - 8(5) + 3(5) = 2 - 40 + 15 = - 40 + 17 = - 23. (Correct answer B.)

OpenStudy (cmtboy2):

i did two thing's i did 2 – 85 + 35=-48 and i did 2 – 8.5 + 3.5=4

OpenStudy (dakid88):

did u look at wht i posted @cmtboy2

OpenStudy (cmtboy2):

ya

Nnesha (nnesha):

it's 8x `8 times x` so you should multiply the 8 with x value put the parentheses

OpenStudy (dakid88):

therefore the answer is B.

OpenStudy (cmtboy2):

no

OpenStudy (dakid88):

no wht

OpenStudy (cmtboy2):

not you @dakid88

Nnesha (nnesha):

so when you replace x for 5 that's how you should write 2-8(5)+3(5) and remember the PEMDAS rule

OpenStudy (dakid88):

oh

OpenStudy (cmtboy2):

oh that i diden't do

OpenStudy (cmtboy2):

Evaluate the expression, using the given value of the variable. z ÷ 4 – 8z + 2(z – 5) when z = 12 A. –79 B. –74 C. –34 D. 107

OpenStudy (cmtboy2):

same here right

OpenStudy (dakid88):

z ÷ 4 – 8z + 2(z – 5) (12)÷4-8(12) +2((12) -5) =-74

Nnesha (nnesha):

stop giving out answer plz

OpenStudy (dakid88):

nope im helping him at the same times

Nnesha (nnesha):

let him/her learn its not ur knowledge test

Nnesha (nnesha):

there is no learning process when u provide the answer as well

OpenStudy (cmtboy2):

im a boy thank you

OpenStudy (dakid88):

lol

OpenStudy (dakid88):

like i said im helping him at the same time @Nnesha thank u

OpenStudy (cmtboy2):

don't you see cmtBOY2

Nnesha (nnesha):

there is a difference between helping and enabling

OpenStudy (dakid88):

leace me alone were most focus on helping @cmtboy2

Nnesha (nnesha):

a girl can use boy word for username

OpenStudy (cmtboy2):

ok but im a boy

OpenStudy (cmtboy2):

how do you do this

OpenStudy (anonymous):

At least try it @cmtboy2, it's the same thing as before. Substitute your variable for your value then simplify.

OpenStudy (anonymous):

hey nnesha not cool dude if the guy is a guy cool if ntot cool just live with it kay:)

Nnesha (nnesha):

did you understand first question ? if you did then you shouldn't ask that question PEMDAS thats what ur teacher wants u to learn

Nnesha (nnesha):

idc if its guy or girl doesn't matter there is a reason i said /him/her its not important to know the gender

OpenStudy (anonymous):

@broski532, you literally added nothing to this discussion. If you're not going to help with the question, don't bother answering.

OpenStudy (cmtboy2):

be nice @tom982

OpenStudy (anonymous):

You're joking, right? @Nnesha was doing a great job helping and @broski532 comes in making an issue out of nothing. We're here to help you guys understand maths, not cause unnecessary drama or just post answers.

OpenStudy (cmtboy2):

am i adding (4)-2 to 12/(4)

Nnesha (nnesha):

nice \[\frac{ 12-4 }{ 12 \div 4 }\]

OpenStudy (cmtboy2):

so i figer them out frist then i add them

OpenStudy (cmtboy2):

2+3

Nnesha (nnesha):

what do you mean *figure them* PEMDAS P) parentheses first E)exponent 3rd MD ) multiply or divide 4th) add or subtract start with the first step is there any parentheses ? no move to the 2nd step is there any exponent ? if no then 3rd step is divide or multiply

Nnesha (nnesha):

nvm there is typo

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha nice \[\frac{ 12-4 }{ 12 \div 4 }\] \(\color{blue}{\text{End of Quote}}\) i meant \[\frac{ -4-2}{ 12 \div- 4 }\] k value is negative

OpenStudy (cmtboy2):

-4-2=6

OpenStudy (cmtboy2):

12/-4=-3

OpenStudy (cmtboy2):

so now we add (-3) to 6

Nnesha (nnesha):

when the sign are the same of both numbers you should just simply add them together and then keep the same sign

Nnesha (nnesha):

-4-2= ?

OpenStudy (cmtboy2):

-6

Nnesha (nnesha):

correct \[\frac{ \color{Red}{-4-2 }}{\color{blue}{ 12 \div -4 }}=\frac{\color{reD}{ -6 }}{ \color{blue}{-3} }\] now just simplify that

OpenStudy (cmtboy2):

now add -6 to -3

Nnesha (nnesha):

it's not -6 `+`(-3) fraction = division

OpenStudy (cmtboy2):

soo divide -3 to -6 right

Nnesha (nnesha):

\[\rm \frac{ \top op }{bottom }\] top part divided by the bottom -3/-6 isn't the same as -6/-3

OpenStudy (cmtboy2):

ok thank

OpenStudy (mathstudent55):

z ÷ 4 – 8z + 2(z – 5) when z = 12 Answer is not -74